OpenStudy (anonymous):

Hi. Is anyone an AP CALCULUS AB GENIUS and interested in helping me with many AP Calculus AB questions? I am falling behind in my class and have in this class. I am learning things like reiman sums. I appreciate your help.

OpenStudy (solomonzelman):

\(\color{#000000}{ \small \displaystyle {\rm AP+CALCULUS+AB+GENIUS}={\rm A\left\{P+C^2(LU)^2S+B\right\}}{\rm +G\in\mathbb{N}} i\cup\int }\)